OpenStudy (anonymous):

a line passes through the point (-3,-4) and its y-intercept is (0,9). what is the equation of the line that is parallel to the first line and passes through the point (3,-7)? a. y=3/5x+2 b. y=5/3x+2 c. y=-5/3x-2 d. y=5/3x-2

Directrix (directrix):

(-3,-4) and its y-intercept is (0,9). This line has slope 13/3. [(9- (-4)]/[0 -(-3)]= 13/3 Parallel lines have the same slope. line that is parallel to the first line and passes through the point (3,-7)? y = (13/3) x - 20 --> Answer (not a given option)

OpenStudy (anonymous):

what is the right answer?

Directrix (directrix):

@2012D > y = (13/3) x - 20 is the correct answer but is not one of the options listed by the OP.

Directrix (directrix):

@2012D If you have a copy of the problem, check to determine if it is posted correctly.

OpenStudy (anonymous):

You have to simplify

OpenStudy (anonymous):

so what's the answer?

OpenStudy (anonymous):

the question was wrong it was supposed to be (0,-9) not (0,9). I still dont understand how to come up with te answer though. :(